calculate the effective nuclear charge on an electron being added to each of tjge following species.
1) F
2) O
3) Ne
4) Cl
Answer #1

The effective nuclear charge calculated with this following equation: Zef=Z-S Where Zell effective nuclear charge Z = atomic4. C1 Cl atomic number = 17. Electronic configuration is 1s 2s 2p 3s 3p The valence cell electrons = 5+2 = 7 electrons. Then